# A Survey on Data Selection for LLM Instruction Tuning

### v2

@@ Abstract.
copyright: ccdoi: 10.1613/jair.1.xxxxxjournalvolume: 4article: 6publicationmonth: 8journalyear: 2025
@@ 1. Introduction
Large language models (LLMs), such as PaLM (Chowdhery et al., 2023), GPT-4 (OpenAI, 2023), and LLaMA (Touvron et al., 2023), have demonstrated impressive capabilities across a wide range of natural language understanding and generation tasks. Central to their success is instruction tuning—a process that aligns model outputs with human intent by fine-tuning pretrained models on curated (instruction, response) pairs (Ouyang et al., 2022b; Zhang et al., 2023). This technique not only improves the controllability and safety of LLMs but also enables them to adapt to specific domains with minimal computational overhead.
In the existing works (Ghosh et al., 2024; Wei et al., 2021; Longpre et al., 2023), instruction tuning mainly focuses on constructing large-scale instruction-following datasets. These works can be categorized into two types: (1) converting annotated NLP datasets into instruction-style formats using templates (e.g., P3 (Sanh et al., 2022) and FLAN (Wei et al., 2021)), and (2) automatically generating instruction-response pairs using powerful models like GPT-3.5 (e.g., Self-Instruct (Wang et al., 2023)). While these works have
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em,font=,, where level=2text width=12em,font=,, where level=3text width=8em,font=,, where level=4text width=5em,font=,, [ Data Selection for Instruction Tuning, ver [Instruction Sets (§2) [ Alpaca (Taori et al., 2023), WizardLM (Xu et al., 2023a), Dolly-V2 (Conover et al., 2023), P3 (Sanh et al., 2022), LIMA (Zhou et al., 2023a), Self-Instruct (Wang et al., 2023) , leaf, text width=36em ] ] [Data Selection Methods (§3) [System of indicators (§3.1) [ INSTRUCTMINING (Cao et al., 2023), DQ (Zhou et al., 2023b),
InstructionGPT-4 (Wei et al., 2023) , leaf, text width=18em ] ] [Trainable LLMs (§3.2) [ IFD (Li et al., 2024c), Instruction Backtranslation (Li et al., 2024b),
Nuggets (Li et al., 2024a), DIVERSEEVOL (Wu et al., 2023),
TEGIT (Chen et al., 2023a), ActiveIT (Kung et al., 2023) , leaf, text width=18em ] ] [Powerful LLMs (§3.3) [ AlpaGasus (Chen et al., 2024), INSTAG (Lu et al., 2024),
LIFT (Xu et al., 2023b), DEITA (Liu et al., 2024),
tree-instruct (Zhao et al., 2023), WaveCoder (Yu et al., 2023) , leaf, text width=18em ] ] [Small Models (§3.4) [ MoDS (Du et al., 2023), Coreset-based Selection (Chen et al., 2023b) , leaf, text width=18em ] ] ] [Evaluation Methods (§4) [ Wining rate (§4.1), Inner Comparison (§4.2), External Comparison (§4.3), leaf, text width=36em ] ] ] ]
@@
In contrast to the previous approaches primarily focused on large-scale data construction (Taori et al., 2023; Xu et al., 2023a; Conover et al., 2023), LIMA (Zhou et al., 2023a) proposes an approach centered on careful data selection. Their findings demonstrated that a model fine-tuned on merely 1,000 carefully curated instruction examples could achieve performance comparable, and in some cases superior, to models trained on vastly larger, often automatically generated, datasets. This seminal work underscored a critical insight: the quality of instruction data can significantly outweigh its quantity in determining the efficacy of instruction tuning. Consequently, this discovery has prompted a shift in research focus, moving from the large-scale generation of instruction data towards developing better methods for constructing and selecting high-quality, impactful, and efficient datasets.
In these data selection approaches, manual instruction data selection often involves high costs and introduces human bias. As a result, automated methods are crucial for instruction data selection. These methods have shown promising results. For instance, the IFD(Li et al., 2024c) method significantly outperforms the Alpaca model by using only about 5% of the Alpaca dataset, and it also surpasses the WizardLM model by about 10%. Using high-quality subsets for fine-tuning not only enhances the instruction-following skills of LLMs but also significantly reduces computational costs and time.
@@ 2. Instruction Datasets
Self-instruct (Wang et al., 2023) consists of 52,000 training and 252 testing instructions. Initial instructions from seed tasks were selected, categorized, and diversified using InstructGPT (Ouyang et al., 2022a) to generate inputs and outputs through output-first or input-first strategies. Post-processing refined the dataset for uniqueness and relevance, providing a versatile resource for natural language processing applications.
Alpaca (Taori et al., 2023) is a dataset of 52,002 instruction-response pairs generated using the Self-Instruct framework (Wang et al., 2023) with the text-davinci-003 model. It is designed to fine-tune LLaMA for improved instruction-following capabilities.
WizardLM (Xu et al., 2023a) is a dataset containing 250,000 instruction-response pairs generated via evolutionary algorithms. Two types of algorithms (depth evolution and breadth evolution) are used to enhance the complexity and scope of basic instructions, generating more complex and diverse high-quality instruction data by ChatGPT.
LIMA (Zhou et al., 2023a) is a dataset consisting of 1,000 training samples, 300 test samples, and 50 development samples. In addition to the manually authored samples, samples collected from Q&A websites are strictly selected by human. Despite its modest scale, LIMA stands out for its meticulous compilation and design. LLMs fine-tuned on LIMA show remarkable ability in following instructions and adapting to unseen tasks.
DollyV2 (Conover et al., 2023) is a dataset of 15,000 human-authored instruction-response pairs covering tasks such as brainstorming, classification, question answering, and summarization. The data was created by employees using only Wikipedia as a source, with restrictions against using web content or generative AI tools when crafting responses.
P3 (Sanh et al., 2022) is a large-scale prompt collection that integrates 170 NLP datasets with 2,052 handcrafted prompts, enabling unified training and evaluation across diverse language tasks. These prompts, also known as task templates, convert conventional NLP tasks like question answering or text classification into natural language input-output pairs. The P3 dataset itself is formed by randomly selecting prompts from PromptSource and organizing data into triplets of inputs, answer choices, and targets.
@@ 3.1. Methods Based on a System of Indicators
INSTRUCTMINING (Cao et al., 2023) is a linear rule-based method for assessing the quality of instructional data. This approach initially identifies key natural language metrics, such as instruction length, perplexity, reward scores, KNN-i(Reimers and Gurevych, 2019; Dong et al., 2011), among others. These metrics are subsequently utilized to develop a linear equation. To explore the correlation between data quality and these metrics, and to define the parameters of equation, comprehensive fine-tuning experiments are conducted. Diverse quality datasets are partitioned into subsets, combined, and then employed in the fine-tuning of large-scale models. Quality labels for each subset are derived by evaluating performance of the model on a test set. The least squares method is applied to these experimental outcomes to estimate the parameters in INSTRUCTMINING. This involves fitting a linear equation to the loss observed when evaluating the model against the test set.Once the parameters have been determined, this formula can be utilized to calculate the quality of instructions, thereby facilitating data selection.
InstructionGPT-4 (Wei et al., 2023) is a data selection method designed for fine-tuning multimodal large language models. It surpasses MiniGPT-4 (Zhu et al., 2024) in various assessments using less data. The first step, metrics like the CLIP Score(Radford et al., 2021), instruction length, and others are used. Visual and textual data are encoded into vectors and then subjected to dimensional reduction, which is regarded as special metrics. These metrics are combined into a vector $e$. The vector is then fed to a trainable data selector, such as a multi-layer perceptron or a self-attention network. This approach, similar to the method (Cao et al., 2023) of calculating quality labels, employs clustering algorithms to segment datasets. Quality labels are assigned post fine-tuning and evaluation of each subset.
DQ (Zhou et al., 2023b) is an innovative data compression approach originally developed for large-scale computer vision datasets, and later adapted for use in the LLM domain. The approach involves several key steps: Initially, a gain function $P(x)$ is defined as:
(5) $P(x)=\sum_{p\in S}||f(p)-f(x)||^{2}_{2}-\sum_{p\in D\setminus S}||f(p)-f(x)||^{2}_{2},$
@@ 3.2. Methods Based on Trainable LLMs
Instruction following difficulty (IFD) (Li et al., 2024c) represents a trainable data selection approach grounded in the hypothesis that large language models (LLMs) can first be trained on a carefully curated subset of instruction data to acquire basic instruction-following capabilities. This foundational skill then enables the model to better estimate instruction difficulty and overall data quality.
@@
Instruction Backtranslation (Li et al., 2024b) is a method for generating and filtering instructions. Starting with a baseline instruction-following model and a web corpus, the model generates instructions for each web document, forming a dataset. Then the model is fine-tuned with seed instructions to gain basic capabilities. And it autonomously scores each instruction, with those exceeding a set threshold forming a high-score subset for further fine-tuning. This iterative process enhances instruction generation and filtering efficiency.
@@
Nuggets (Li et al., 2024a) is a framework that implements a dual-phase approach. Initially, it employs a variety of predefined tasks to evaluate the proficiency of a LLM across multiple scenarios, a process known as zero-shot scoring. Subsequently, each entry in the instruction dataset is utilized as a unique prompt in a one-shot manner. These prompts are presented before the predefined tasks, and the performance of LLM is reassessed, a step termed one-shot scoring. This approach harnesses the disparity between one-shot and zero-shot scores to calculate a definitive ’gold score’ for each instruction. Upon acquiring gold scores for all instructions, those constituting the highest-scoring subset are chosen as the ’gold subset’. This subset is then directly employed in the fine-tuning of the model. This methodology exploits the inherent contextual learning capabilities of extensive models.
@@
DIVERSEEVOL (Wu et al., 2023) is an innovative iterative data selection mechanism. It leverages large-scale models like LLaMa to generate embedding vectors for instructional data. The mechanism employs the k-center-greedy algorithm(Sener and Savarese, 2018a) to facilitate diversity in selecting a subset of data for fine-tuning the LLaMa model. This procedure is repetitively applied, progressively enlarging the chosen subset, culminating in the creation of a high-quality instructional dataset.
@@
TEGIT (Chen et al., 2023a) is a novel method for generating high-quality instruction fine-tuning data. Particularly notable is their method of filtering instruction data. Utilizing ChatGPT, a small corpus of documentation is transformed into a format suitable for instruction data, forming a meta-dataset. This dataset then serves to train two Llama2 models - one functioning as a task generator and the other as a task discriminator. The role of generator is to devise tasks from provided texts, while the discriminator evaluates these tasks, ensuring their quality.
@@
Active Instruction Tuning (Kung et al., 2023) is a method that focuses on task sensitivity selection to enhance large model fine-tuning with fewer tasks while improving out-of-task generalization. This method introduces the concept of Prompt Uncertainty, which is determined by generating k perturbed instructions through the random deletion of words from an original instruction. The output of LLM probability deviation is then averaged for these k perturbed instructions. Tasks that exhibit higher Prompt Uncertainty are prioritized for instruction fine-tuning, using the extent of Prompt Uncertainty as a measure of task uncertainty.
@@ 3.3. Methods based on Powerful LLMs like ChatGPT
ALPAGASUS (Chen et al., 2024) is an innovative data filtering methodology aimed at improving efficiency and accuracy in Instruction-Following Task (IFT) data curation. This approach utilizes a well-designed prompt applied to ChatGPT, to assess the quality of each data tuple, comprising instruction, input, and response. The methodology focuses on excluding tuples that fall below a predefined quality threshold. When this filtering process was applied to a substantial dataset, it was observed that a considerable portion of the data was compromised by quality issues. Notably, the application of the LLM-based filtering process led to the development of a model that surpassed the performance of the original model, which was trained using the unfiltered dataset with instruction-based fine-tuning.
@@
INSTAG (Lu et al., 2024) is an automated instruction tagging method that employs ChatGPT to generate detailed, open-ended labels for instructions, enhancing the diversity and complexity of the resulting subset. The process involves annotating data with labels that reflect the semantic and intent of each instruction and normalizing them for a label-based selection approach. This approach follows a complexity-first diversification sampling strategy. It starts by sorting queries in descending order of label count and then iteratively adds queries to a subset based on the uniqueness of their labels, until the subset reaches a desired size N. The result is a curated set of queries with a higher number of labels, indicating greater complexity and diversity.
@@
To enhance the distribution and quality of datasets, (Xu et al., 2023b) propose the method LIFT to reduce the redundancy of samples. It involves two stages: expanding dataset distribution and curating the diversity and quality of the dataset. Initially, ChatGPT enhances the data by generating diverse instructions and vectorizing them. A subset is then selected based on row variances. Secondly, ChatGPT scores instructions for accuracy, interpretability, clarity, difficulty, and length. The initial subset is re-selected based on these scores.
@@
DEITA (Liu et al., 2024) integrates a multifaceted approach for selecting instruction data, focusing on complexity, quality, and diversity. Complexity describes factors such as the length, difficulty, and complexity of the instruction. And quality captures the accuracy of the output. Utilizing the WizardLM technique, ChatGPT is employed to augment instructions, which are then evaluated for complexity and quality. These evaluations involve scoring instructions based on complexity, using a specially trained complexity scorer, and assessing the output quality. Each instruction in the dataset is assigned complexity scores ($c$) and quality scores ($q$), and a comprehensive score is calculated by multiplying these two metrics. The dataset is then organized according to these comprehensive scores and vectorized for further analysis. To ensure diversity, a subset is created by adding samples that exceed a set distance threshold ($\tau$) from their nearest neighbor in the subset. This process continues until the subset reaches a predetermined size.
@@
Zhao et al. (2023) propose the tree-instruct method, which improves the quality of instruction by enhance the complexity of it. Liu et al. (2024) uses this method to measure instruction complexity and sets thresholds for instruction filtering. Tree-instruct employs GPT-4 to generate semantic parse trees for instruction data, using the number of nodes in the tree as a measure of complexity. The complexity is enhanced by adding nodes to the tree, and then GPT-4 is used to convert the new tree back into natural language, resulting in new high-quality instructions.
@@
WaveCoder (Yu et al., 2023) is a code-focused LLM enhanced by an instruction improvement technique. Its training incorporates generated data, with the data filtering phase being particularly significant. Following data generation, a LLM-based discriminator, leveraging GPT-4, evaluates instruction data against established criteria divided into subtopics. This approach enables finer control over the filtering process, effectively eliminating low-quality instruction instances.

### v3

@@ Abstract.
copyright: ccdoi: 10.1613/jair.1.17625article: 32journalvolume: 83publicationmonth: 8journalyear: 2025
@@ 1. Introduction
Large language models (LLMs), such as PaLM (PaLM, 5), GPT-4 (gpt4, 17), and LLaMA (LLaMa, 26), have demonstrated impressive capabilities across a wide range of natural language understanding and generation tasks. Central to their success is instruction tuning—a process that aligns model outputs with human intent by fine-tuning pretrained models on curated (instruction, response) pairs (LLM_hf, 18, 34). This technique not only improves the controllability and safety of LLMs but also enables them to adapt to specific domains with minimal computational overhead.
In the existing works (closer, 9, 28, 15), instruction tuning mainly focuses on constructing large-scale instruction-following datasets. These works can be categorized into two types: (1) converting annotated NLP datasets into instruction-style formats using templates (e.g., P3 (P3, 22) and FLAN (FLAN, 28)), and (2) automatically generating instruction-response pairs using powerful models like GPT-3.5 (e.g., Self-Instruct (Self-Instruct, 27)). While these works have
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=12em,font=,, where level=2text width=12em,font=,, where level=3text width=8em,font=,, where level=4text width=5em,font=,, [ Data Selection for Instruction Tuning, ver [Instruction Sets (§2) [ Alpaca (alpaca, 25), WizardLM (WizardLM, 31), Dolly-V2 (DollyV2, 6), P3 (P3, 22), LIMA (zhou2023lima, 36), Self-Instruct (Self-Instruct, 27) , leaf, text width=36em ] ] [Data Selection Methods (§3) [System of indicators (§3.1) [ INSTRUCTMINING (InstructionMining, 1), DQ (DatasetQuantization, 37),
InstructionGPT-4 (InstructionGPT-4, 29) , leaf, text width=18em ] ] [Trainable LLMs (§3.2) [ IFD (IFD, 11), Instruction Backtranslation (Instruction_Backtranslation, 12),
Nuggets (One_Shot_Learning, 13), DIVERSEEVOL (Self-Evolved, 30),
TEGIT (TeGit, 4), ActiveIT (ActiveIT, 10) , leaf, text width=18em ] ] [Powerful LLMs (§3.3) [ AlpaGasus (AlpaGasus, 3), INSTAG (InsTag, 16),
LIFT (LIFT, 32), DEITA (DEITA, 14),
InstructionNode (InstructionNode, 35), WaveCoder (WaveCoder, 33) , leaf, text width=18em ] ] [Small Models (§3.4) [ MoDS (MoDS, 8), Coreset-based Selection (0.5data, 2) , leaf, text width=18em ] ] ] [Evaluation Methods (§4) [ Wining rate (§4.1), Inner Comparison (§4.2), External Comparison (§4.3), leaf, text width=36em ] ] ] ]
@@
In contrast to the previous approaches primarily focused on large-scale data construction (alpaca, 25, 31, 6), LIMA (zhou2023lima, 36) proposes an approach centered on careful data selection. Their findings demonstrated that a model fine-tuned on merely 1,000 carefully curated instruction examples could achieve performance comparable, and in some cases superior, to models trained on vastly larger, often automatically generated, datasets. This seminal work underscored a critical insight: the quality of instruction data can significantly outweigh its quantity in determining the efficacy of instruction tuning. Consequently, this discovery has prompted a shift in research focus, moving from the large-scale generation of instruction data towards developing better methods for constructing and selecting high-quality, impactful, and efficient datasets.
In these data selection approaches, manual instruction data selection often involves high costs and introduces human bias. As a result, automated methods are crucial for instruction data selection. These methods have shown promising results. For instance, the IFD(IFD, 11) method significantly outperforms the Alpaca model by using only about 5% of the Alpaca dataset, and it also surpasses the WizardLM model by about 10%. Using high-quality subsets for fine-tuning not only enhances the instruction-following skills of LLMs but also significantly reduces computational costs and time.
@@ 2. Instruction Datasets
Self-instruct (Self-Instruct, 27) consists of 52,000 training and 252 testing instructions. Initial instructions from seed tasks were selected, categorized, and diversified using InstructGPT (Instructgpt, 19) to generate inputs and outputs through output-first or input-first strategies. Post-processing refined the dataset for uniqueness and relevance, providing a versatile resource for natural language processing applications.
Alpaca (alpaca, 25) is a dataset of 52,002 instruction-response pairs generated using the Self-Instruct framework (Self-Instruct, 27) with the text-davinci-003 model. It is designed to fine-tune LLaMA for improved instruction-following capabilities.
WizardLM (WizardLM, 31) is a dataset containing 250,000 instruction-response pairs generated via evolutionary algorithms. Two types of algorithms (depth evolution and breadth evolution) are used to enhance the complexity and scope of basic instructions, generating more complex and diverse high-quality instruction data by ChatGPT.
LIMA (zhou2023lima, 36) is a dataset consisting of 1,000 training samples, 300 test samples, and 50 development samples. In addition to the manually authored samples, samples collected from Q&A websites are strictly selected by human. Despite its modest scale, LIMA stands out for its meticulous compilation and design. LLMs fine-tuned on LIMA show remarkable ability in following instructions and adapting to unseen tasks.
DollyV2 (DollyV2, 6) is a dataset of 15,000 human-authored instruction-response pairs covering tasks such as brainstorming, classification, question answering, and summarization. The data was created by employees using only Wikipedia as a source, with restrictions against using web content or generative AI tools when crafting responses.
P3 (P3, 22) is a large-scale prompt collection that integrates 170 NLP datasets with 2,052 handcrafted prompts, enabling unified training and evaluation across diverse language tasks. These prompts, also known as task templates, convert conventional NLP tasks like question answering or text classification into natural language input-output pairs. The P3 dataset itself is formed by randomly selecting prompts from PromptSource and organizing data into triplets of inputs, answer choices, and targets.
@@ 3.1. Methods Based on a System of Indicators
INSTRUCTMINING (InstructionMining, 1) is a linear rule-based method for assessing the quality of instructional data. This approach initially identifies key natural language metrics, such as instruction length, perplexity, reward scores, KNN-i(Sentence-BERT, 21, 7), among others. These metrics are subsequently utilized to develop a linear equation. To explore the correlation between data quality and these metrics, and to define the parameters of equation, comprehensive fine-tuning experiments are conducted. Diverse quality datasets are partitioned into subsets, combined, and then employed in the fine-tuning of large-scale models. Quality labels for each subset are derived by evaluating performance of the model on a test set. The least squares method is applied to these experimental outcomes to estimate the parameters in INSTRUCTMINING. This involves fitting a linear equation to the loss observed when evaluating the model against the test set.Once the parameters have been determined, this formula can be utilized to calculate the quality of instructions, thereby facilitating data selection.
InstructionGPT-4 (InstructionGPT-4, 29) is a data selection method designed for fine-tuning multimodal large language models. It surpasses MiniGPT-4 (MiniGPT-4, 38) in various assessments using less data. The first step, metrics like the CLIP Score(CLIP, 20), instruction length, and others are used. Visual and textual data are encoded into vectors and then subjected to dimensional reduction, which is regarded as special metrics. These metrics are combined into a vector $e$. The vector is then fed to a trainable data selector, such as a multi-layer perceptron or a self-attention network. This approach, similar to the method (InstructionMining, 1) of calculating quality labels, employs clustering algorithms to segment datasets. Quality labels are assigned post fine-tuning and evaluation of each subset.
DQ (DatasetQuantization, 37) is an innovative data compression approach originally developed for large-scale computer vision datasets, and later adapted for use in the LLM domain. The approach involves several key steps: Initially, a gain function $P(x)$ is defined as:
[…]
@@ 3.2. Methods Based on Trainable LLMs
Instruction following difficulty (IFD) (IFD, 11) represents a trainable data selection approach grounded in the hypothesis that large language models (LLMs) can first be trained on a carefully curated subset of instruction data to acquire basic instruction-following capabilities. This foundational skill then enables the model to better estimate instruction difficulty and overall data quality.
@@
Instruction Backtranslation (Instruction_Backtranslation, 12) is a method for generating and filtering instructions. Starting with a baseline instruction-following model and a web corpus, the model generates instructions for each web document, forming a dataset. Then the model is fine-tuned with seed instructions to gain basic capabilities. And it autonomously scores each instruction, with those exceeding a set threshold forming a high-score subset for further fine-tuning. This iterative process enhances instruction generation and filtering efficiency.
@@
Nuggets (One_Shot_Learning, 13) is a framework that implements a dual-phase approach. Initially, it employs a variety of predefined tasks to evaluate the proficiency of a LLM across multiple scenarios, a process known as zero-shot scoring. Subsequently, each entry in the instruction dataset is utilized as a unique prompt in a one-shot manner. These prompts are presented before the predefined tasks, and the performance of LLM is reassessed, a step termed one-shot scoring. This approach harnesses the disparity between one-shot and zero-shot scores to calculate a definitive ’gold score’ for each instruction. Upon acquiring gold scores for all instructions, those constituting the highest-scoring subset are chosen as the ’gold subset’. This subset is then directly employed in the fine-tuning of the model. This methodology exploits the inherent contextual learning capabilities of extensive models.
@@
DIVERSEEVOL (Self-Evolved, 30) is an innovative iterative data selection mechanism. It leverages large-scale models like LLaMa to generate embedding vectors for instructional data. The mechanism employs the k-center-greedy algorithm(k-center-greedy, 23) to facilitate diversity in selecting a subset of data for fine-tuning the LLaMa model. This procedure is repetitively applied, progressively enlarging the chosen subset, culminating in the creation of a high-quality instructional dataset.
@@
TEGIT (TeGit, 4) is a novel method for generating high-quality instruction fine-tuning data. Particularly notable is their method of filtering instruction data. Utilizing ChatGPT, a small corpus of documentation is transformed into a format suitable for instruction data, forming a meta-dataset. This dataset then serves to train two Llama2 models - one functioning as a task generator and the other as a task discriminator. The role of generator is to devise tasks from provided texts, while the discriminator evaluates these tasks, ensuring their quality.
@@
Active Instruction Tuning (ActiveIT, 10) is a method that focuses on task sensitivity selection to enhance large model fine-tuning with fewer tasks while improving out-of-task generalization. This method introduces the concept of Prompt Uncertainty, which is determined by generating k perturbed instructions through the random deletion of words from an original instruction. The output of LLM probability deviation is then averaged for these k perturbed instructions. Tasks that exhibit higher Prompt Uncertainty are prioritized for instruction fine-tuning, using the extent of Prompt Uncertainty as a measure of task uncertainty.
@@ 3.3. Methods Based on Powerful LLMs like ChatGPT
ALPAGASUS (AlpaGasus, 3) is an innovative data filtering methodology aimed at improving efficiency and accuracy in Instruction-Following Task (IFT) data curation. This approach utilizes a well-designed prompt applied to ChatGPT, to assess the quality of each data tuple, comprising instruction, input, and response. The methodology focuses on excluding tuples that fall below a predefined quality threshold. When this filtering process was applied to a substantial dataset, it was observed that a considerable portion of the data was compromised by quality issues. Notably, the application of the LLM-based filtering process led to the development of a model that surpassed the performance of the original model, which was trained using the unfiltered dataset with instruction-based fine-tuning.
@@
INSTAG (InsTag, 16) is an automated instruction tagging method that employs ChatGPT to generate detailed, open-ended labels for instructions, enhancing the diversity and complexity of the resulting subset. The process involves annotating data with labels that reflect the semantic and intent of each instruction and normalizing them for a label-based selection approach. This approach follows a complexity-first diversification sampling strategy. It starts by sorting queries in descending order of label count and then iteratively adds queries to a subset based on the uniqueness of their labels, until the subset reaches a desired size N. The result is a curated set of queries with a higher number of labels, indicating greater complexity and diversity.
@@
To enhance the distribution and quality of datasets, (LIFT, 32) propose the method LIFT to reduce the redundancy of samples. It involves two stages: expanding dataset distribution and curating the diversity and quality of the dataset. Initially, ChatGPT enhances the data by generating diverse instructions and vectorizing them. A subset is then selected based on row variances. Secondly, ChatGPT scores instructions for accuracy, interpretability, clarity, difficulty, and length. The initial subset is re-selected based on these scores.
@@
DEITA (DEITA, 14) integrates a multifaceted approach for selecting instruction data, focusing on complexity, quality, and diversity. Complexity describes factors such as the length, difficulty, and complexity of the instruction. And quality captures the accuracy of the output. Utilizing the WizardLM technique, ChatGPT is employed to augment instructions, which are then evaluated for complexity and quality. These evaluations involve scoring instructions based on complexity, using a specially trained complexity scorer, and assessing the output quality. Each instruction in the dataset is assigned complexity scores ($c$) and quality scores ($q$), and a comprehensive score is calculated by multiplying these two metrics. The dataset is then organized according to these comprehensive scores and vectorized for further analysis. To ensure diversity, a subset is created by adding samples that exceed a set distance threshold ($\tau$) from their nearest neighbor in the subset. This process continues until the subset reaches a predetermined size.
@@
InstructionNode (InstructionNode, 35) is a tree-instruct method, which improves the quality of instruction by enhance the complexity of it. DEITA (DEITA, 14) uses this method to measure instruction complexity and sets thresholds for instruction filtering. Tree-instruct employs GPT-4 to generate semantic parse trees for instruction data, using the number of nodes in the tree as a measure of complexity. The complexity is enhanced by adding nodes to the tree, and then GPT-4 is used to convert the new tree back into natural language, resulting in new high-quality instructions.
@@
WaveCoder (WaveCoder, 33) is a code-focused LLM enhanced by an instruction improvement technique. Its training incorporates generated data, with the data filtering phase being particularly significant. Following data generation, a LLM-based discriminator, leveraging GPT-4, evaluates instruction data against established criteria divided into subtopics. This approach enables finer control over the filtering process, effectively eliminating low-quality instruction instances.